# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
    Assuming instead that $S_{F}\subseteq S_{\xi}$, then ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective if and only if there exists an irreducible component ${\mathcal{H}}_{F^{\prime}}$ of ${\mathcal{H}}_{F}$ such that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)$ is not equal to a twisted sign or twisted trivial character. Equivalently, this holds if and only if there exists $s,s^{\prime}\in S_{F}$ corresponding to adjacent nodes in the affine Dynkin diagram such that $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$.
@@
To prove non-projectivity, it suffices to find $F\subseteq\overline{F^{\prime\prime}}\subseteq\overline{C}$ such that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective, because ${\mathcal{H}}_{F}$ is free over ${\mathcal{H}}_{F^{\prime\prime}}$. For (i), we pick $F^{\prime\prime}$ to be given by $S_{F^{\prime\prime}}=\{s\}$, where $s\in S_{F}\setminus S_{\xi}$. That ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective is then the content of the proof of [19, Lemma 7.5]. For (ii), we consider the idempotent
$e_{\xi}=\left|T(\mathbb{F}_{q})\right|^{-1}\sum_{t\in T(\mathbb{F}_{q})}\xi(t)T_{t^{-1}}\in k[T(\mathbb{F}_{q})]$
corresponding to $\xi$, and note that $e_{\xi}$ is central in ${\mathcal{H}}_{F}$ since $S_{F}\subseteq S_{\xi}$ by assumption. It follows that the character ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ factors through the product of blocks $e_{\xi}{\mathcal{H}}_{F}$. This algebra is the 00-Hecke algebra associated to the finite subgroup $W_{F}\leq W$ generated by $S_{F}$, and is the tensor product of the 00-Hecke algebras $e_{\xi}{\mathcal{H}}_{F^{\prime}}$ corresponding to the irreducible components of ${\mathcal{H}}_{F}$. But it is shown in [23, Lemma 4.2 & Theorem 5.2] that the trivial and sign are the only projective characters over the 0-Hecke algebra of an irreducible root system. Thus, if the restriction of $\chi$ to each irreducible component of ${\mathcal{H}}_{F}$ is a twisted sign or trivial character we obtain that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F}}(\chi)$ is projective, and then that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective as well since $e_{\xi}{\mathcal{H}}_{F}$ is a product of blocks of ${\mathcal{H}}_{F}$.
@@
Conversely, suppose that there exists $s,s^{\prime}\in S_{F}$ adjacent in the affine Dynkin diagram with $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$. Then set $S_{F^{\prime\prime}}=\{s,s^{\prime}\}$, and once again we have that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ factors through the product of blocks $e_{\xi}{\mathcal{H}}_{F^{\prime\prime}}$. The latter is isomorphic to a 0-Hecke algebra of type $A_{2},$ $B_{2}$ or $G_{2}$, and hence it follows from loc. cit. that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective, since it is neither a sign nor a trivial character by the condition that $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$. As $e_{\xi}{\mathcal{H}}_{F^{\prime\prime}}$ is a product of blocks of ${\mathcal{H}}_{F^{\prime\prime}}$, we are done. ∎
@@
We now make a couple of preliminary observations. Suppose that we have a collection of morphisms as in 5.9 and fix a face $F\subseteq\overline{C}$. If ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective then $f_{F}\sim 0$. Similarly, if there is a face $F^{\prime}\subseteq\overline{F}$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ then $f_{F}\sim 0$ by 5.5(i) and 5.9(i). We therefore only need to consider a face $F$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective and the restrictions of $\chi$ and $\chi^{\prime}$ to ${\mathcal{H}}_{F^{\prime}}$ agree for all $F^{\prime}\subseteq\overline{F}$. In particular, note that the condition ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi^{\prime})$ implies that $\xi=\xi^{\prime}$ since $k[T(\mathbb{F}_{q})]={\mathcal{H}}_{C}\subseteq{\mathcal{H}}_{F}$.
Assume from now on that we are in the above situation. Then the set
@@
must be non-empty since $\chi\neq\chi^{\prime}$ and $\xi=\xi^{\prime}$, and we denote its elements by $s_{1},\ldots,s_{n}$. For each $1\leq j\leq n$, we then write $S_{i_{j}}$ for the irreducible component containing $s_{j}$. Observe further that our assumptions on $F$, $\chi$ and $\chi^{\prime}$ imply that for each $j$ there is no face $F^{\prime}\subseteq\overline{F}$ with $S_{F^{\prime}}=S_{F}\cup\{s_{j}\}$. In other words, we must have $S_{F}\cap S_{i_{j}}=S_{i_{j}}\setminus\{s_{j}\}$ for every $1\leq j\leq n$ by (2.3). In particular, the corresponding components $\Phi_{i_{1}},\ldots,\Phi_{i_{n}}$ are all distinct.
@@
Suppose $\chi\neq\chi^{\prime}$ are supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ and that $\underline{f}:=(f_{F})\in\prod_{F\subseteq\overline{C}}{\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$ is as in 5.9. Assume $F\subseteq\overline{C}$ is a face such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective and ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ whenever $F^{\prime}\subseteq\overline{F}$. With the above notation, we have:
@@
(i) As explained above we must have $\xi=\xi^{\prime}$ and $S_{F}\cap S_{i_{j}}=S_{i_{j}}\setminus\{s_{j}\}$ for all $1\leq j\leq n$. Suppose $\Phi_{i_{j}}$ is of type $A_{1}$ for some $j$. Then $S_{i_{j}}$ consists of two simple affine reflection $s_{j}$ and $s^{\prime}_{j}$, say, with the same associated coroot. Since $\chi(T_{\hat{s_{j}}})\neq\chi^{\prime}(T_{\hat{s_{j}}})$ we must have $s_{j}\in S_{\xi}$ and so $S_{\xi}\cap S_{i_{j}}=S_{i_{j}}$ by definition of $S_{\xi}$. Without loss of generality, we assume $\chi(T_{\hat{s_{j}}})=-1$ and $\chi^{\prime}(T_{\hat{s_{j}}})=0$. Then by supersingularity of $\chi$ and $\chi^{\prime}$ we must have $\chi(T_{\hat{s^{\prime}_{j}}})=0$ and $\chi^{\prime}(T_{\hat{s^{\prime}_{j}}})=-1$. But we also have that $s^{\prime}_{j}\in S_{F}$ so that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi^{\prime})$, contradicting our assumptions.
(ii) Let $s^{\prime}\in S_{F}\setminus S_{\xi}$ and note as above that we necessarily have $s^{\prime}\neq s_{j}$ for any $1\leq j\leq n$. We can now define $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $F=F_{1}$, $S_{F_{2}}=\{s_{1},s^{\prime}\}$ and $S_{F_{3}}=\{s^{\prime}\}$. Note that such a face $F_{2}$ exists because of (i). 5.10(i) then tells us that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective. So 5.9(iii) gives us that $f_{F}\sim 0$ as required. ∎
@@
Suppose that $\chi$, $\chi^{\prime}$, $\underline{f}$ and $F\subseteq\overline{C}$ are given as in 5.11, and assume further that $S_{F}\subseteq S_{\xi}$.
@@
      $s^{\prime},s^{\prime\prime}\notin S_{i_{j}}$ for any $1\leq j\leq n$;
@@
The existence of $s^{\prime},s^{\prime\prime}\in S_{F}$ as in (i) follows from 5.10(ii) because ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective by assumption. For (ii), we let $F_{1},F_{2}\subseteq\overline{F_{3}}$ with $F=F_{1}$, $S_{F_{2}}=\{s_{j},s^{\prime},s^{\prime\prime}\}$ and $S_{F_{3}}=\{s^{\prime},s^{\prime\prime}\}$, for a suitably chosen $1\leq j\leq n$ so that such an $F_{2}$ exists. In case (a), any choice of $j$ works. In case (b), we pick $j$ such that $\left|S_{i_{j}}\right|>3$. In case (c), we pick $j$ such that $s^{\prime},s^{\prime\prime}\notin S_{i_{j}}$. Either way, as ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi^{\prime})$, it follows from 5.5(i) and 5.9(iii) that we only need to show that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective, which itself holds by 5.10(ii). ∎
By 5.12(ii) we may simplify our notation a bit. We write $\Phi=\sqcup_{i=0}^{r}\Phi_{i}$ as before and assume that $S_{1}=\{s,s^{\prime},s^{\prime\prime}\}\subseteq S_{\xi}$, where $\chi(T_{\hat{s^{\prime}}})\neq\chi(T_{\hat{s^{\prime\prime}}})$. Furthermore, we assume that $s\in S_{1}$ is the unique simple affine reflection for which $\chi(T_{\hat{s}})\neq\chi^{\prime}(T_{\hat{s}})$.
@@
Let $\chi$, $\chi^{\prime}$, $\underline{f}$ be given as in 5.12, and suppose that $\Phi$, $\chi$ and $\chi^{\prime}$ satisfy the assumptions above. Let $F\subseteq\overline{C}$ be any face with $S_{F}\cap S_{1}=\{s^{\prime},s^{\prime\prime}\}$.

### v4

@@ 5.2. Supersingular characters in ${\operatorname{Ho}}({\mathcal{H}_{\operatorname{aff}}})$
    Assuming instead that $S_{F}\subseteq S_{\xi}$, then the following are equivalent:
    1. (a)
      ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective;
    2. (b)
      there exists an irreducible component ${\mathcal{H}}_{F^{\prime}}$ of ${\mathcal{H}}_{F}$ such that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)$ is not equal to a twisted sign or twisted trivial character; and
    3. (c)
      there exist $s,s^{\prime}\in S_{F}$ corresponding to adjacent nodes in the affine Dynkin diagram such that $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$.
@@
To prove non-projectivity, it suffices to find $F\subseteq\overline{F^{\prime\prime}}\subseteq\overline{C}$ such that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective, because ${\mathcal{H}}_{F}$ is free over ${\mathcal{H}}_{F^{\prime\prime}}$. For (i), we pick $F^{\prime\prime}$ to be given by $S_{F^{\prime\prime}}=\{s\}$, where $s\in S_{F}\setminus S_{\xi}$. That ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective is then the content of the proof of [21, Lemma 7.5]. For (ii), we consider the idempotent $e_{\xi}$ (cf. (5.5)). First note that $e_{\xi}$ is central in ${\mathcal{H}}_{F}$ and that the algebra $e_{\xi}{\mathcal{H}}_{F}$ is the 00-Hecke algebra associated to the finite subgroup $W_{F}\leq W$ generated by $S_{F}$, cf. 5.5. Also, from ${\mathcal{H}}_{F}=e_{\xi}{\mathcal{H}}_{F}\times(1-e_{\xi}){\mathcal{H}}_{F}$ we see that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ factors through the algebra $e_{\xi}{\mathcal{H}}_{F}$. Hence, to show that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective is equivalent to showing that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F}}(\chi)$ is projective. By the braid relations, $e_{\xi}{\mathcal{H}}_{F}$ is the tensor product of the 00-Hecke algebras $e_{\xi}{\mathcal{H}}_{F^{\prime}}$ corresponding to the irreducible components of ${\mathcal{H}}_{F}$, and ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F}}(\chi)$ decomposes similarly as the tensor product of its restrictions to each component.
Now, suppose $\mathcal{C}$ is the 0-Hecke algebra of an irreducible Coxeter system $(W_{\text{gen}},S_{\text{gen}})$. For a character $\lambda_{L}$ as above, we denote by $P_{L}$ its indecomposable projective cover. It is shown in [25, Theorem 5.1 & Theorem 5.2] that the trivial and sign characters of $\mathcal{C}$ are projective, and that for any $L\neq\emptyset$, $S_{\text{gen}}$ there is $L^{\prime}\neq L$ such that $\lambda_{L}$ is a composition factor of $P_{L^{\prime}}$. In particular, $\lambda_{L}$ is never projective for $L\neq\emptyset$, $S_{\text{gen}}$. Going back to our situation, if the restriction of $\chi$ to each irreducible component of ${\mathcal{H}}_{F}$ is a twisted sign or trivial character it then follows that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F}}(\chi)$ is projective, since it is a tensor product of projectives. This shows (a)$\Rightarrow$(b). Moreover, (b)$\Rightarrow$(c) is clear by definition.
@@
Finally, suppose that there exist $s,s^{\prime}\in S_{F}$ adjacent in the affine Dynkin diagram with $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$. Then define $F^{\prime\prime}$ by $S_{F^{\prime\prime}}=\{s,s^{\prime}\}$, and once again note that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ factors through the algebra $e_{\xi}{\mathcal{H}}_{F^{\prime\prime}}$. The latter is isomorphic to a 0-Hecke algebra of type $A_{2},$ $B_{2}$ or $G_{2}$, and hence it follows from loc. cit. that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{e_{\xi}{\mathcal{H}}_{F^{\prime\prime}}}(\chi)$ is not projective, since it is neither a sign nor a trivial character by assumption. This gives (c)$\Rightarrow$(a). ∎
@@
Suppose that we have a collection of morphisms as in 5.11 and fix a face $F\subseteq\overline{C}$. Our aim is to show that unless we are in the setting of 5.8(i) and $F$ is defined by $S_{F}=\{s^{\prime},s^{\prime\prime}\}$ (with $s^{\prime},s^{\prime\prime}$ as in the statement of the Proposition), then $f_{F}\sim 0$. We will do this in a sequence of reductions where we will impose more conditions on $\chi$, $\chi^{\prime}$, $F$ and $\Phi$ as we go, and show that $f_{F}\sim 0$ whenever those conditions fail. For this, first note that if ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is projective then $f_{F}\sim 0$. Similarly, if there is a face $F^{\prime}\subseteq\overline{F}$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ then $f_{F}\sim 0$ by 5.6(i) and 5.11(i). To compute $(f_{F})_{F\subseteq\overline{C}}$, we are thus left to consider a face $F$ such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective and ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ for all $F^{\prime}\subseteq\overline{F}$. In particular, note that this last condition implies that $\xi=\xi^{\prime}$ since $k[T(\mathbb{F}_{q})]={\mathcal{H}}_{C}\subseteq{\mathcal{H}}_{F}$.
Assume from now on that $\xi=\xi^{\prime}$ and that $F$ is as above. Then the set
@@
must be non-empty since $\chi\neq\chi^{\prime}$ and $\xi=\xi^{\prime}$, and we denote its elements by $s_{1},\ldots,s_{n}$. For each $1\leq j\leq n$, we then write $S_{i_{j}}$ for the irreducible component containing $s_{j}$. Observe further that our assumptions on $F$, $\chi$ and $\chi^{\prime}$ imply that for each $j$ there is no face $F^{\prime}\subseteq\overline{F}$ with $S_{F^{\prime}}=S_{F}\cup\{s_{j}\}$. In other words, we must have $S_{F}\cap S_{i_{j}}=S_{i_{j}}\setminus\{s_{j}\}$ for every $1\leq j\leq n$ by (2.3). In particular, the corresponding components $S_{i_{1}},\ldots,S_{i_{n}}$ are all distinct.
We now begin our proof of 5.8 by first showing that $f_{F}\sim 0$ whenever $S_{F}\not\subseteq S_{\xi}$.
@@
Suppose $\chi\neq\chi^{\prime}$ are supersingular characters of ${\mathcal{H}_{\operatorname{aff}}}$ and that $\underline{f}:=(f_{F})\in\prod_{F\subseteq\overline{C}}{\operatorname{Hom}}_{{\mathcal{H}}_{F}}(\chi,\chi^{\prime})$ is as in 5.11. Assume $F\subseteq\overline{C}$ is a face such that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective and ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi)={\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F^{\prime}}}(\chi^{\prime})$ whenever $F^{\prime}\subseteq\overline{F}$. With the above notation, we have:
@@
(i) As explained above we have $S_{F}\cap S_{i_{j}}=S_{i_{j}}\setminus\{s_{j}\}$ for all $1\leq j\leq n$. Suppose $\Phi_{i_{j}}$ is of type $A_{1}$ for some $j$. Then $S_{i_{j}}$ consists of two simple affine reflections $s_{j}$ and $s^{\prime}_{j}$, say, with the same associated coroot. Since $\chi(T_{\hat{s_{j}}})\neq\chi^{\prime}(T_{\hat{s_{j}}})$ we must have $s_{j}\in S_{\xi}$ and so $S_{i_{j}}\subseteq S_{\xi}$ by definition of $S_{\xi}$. Without loss of generality, we assume $\chi(T_{\hat{s_{j}}})=-1$ and $\chi^{\prime}(T_{\hat{s_{j}}})=0$. Then by supersingularity of $\chi$ and $\chi^{\prime}$ we must have $\chi(T_{\hat{s^{\prime}_{j}}})=0$ and $\chi^{\prime}(T_{\hat{s^{\prime}_{j}}})=-1$. But we also have that $s^{\prime}_{j}\in S_{F}$ so that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi^{\prime})$, contradicting our assumptions.
(ii) Let $s^{\prime}\in S_{F}\setminus S_{\xi}$ and note as above that we necessarily have $s^{\prime}\neq s_{j}$ for any $1\leq j\leq n$. We can now define $F_{1},F_{2}\subseteq\overline{F_{3}}$ by $F=F_{1}$, $S_{F_{2}}=\{s_{1},s^{\prime}\}$ and $S_{F_{3}}=\{s^{\prime}\}$. Note that such a face $F_{2}$ exists because of (i). 5.12(i) then tells us that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective. As ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi^{\prime})$, 5.11(iii) gives us that $f_{F}\sim 0$ as required. ∎
Next, in the case where instead $S_{F}\subseteq S_{\xi}$, we reduce quite thoroughly the possibilities for $\chi$ and $\Phi$ in order to avoid $f_{F}\sim 0$.
@@
Suppose that $\chi$, $\chi^{\prime}$, $\underline{f}$ and $F\subseteq\overline{C}$ are given as in 5.13, and assume further that $S_{F}\subseteq S_{\xi}$.
@@
      there are $s^{\prime},s^{\prime\prime}\in S_{F}$ as in (i) such that $s^{\prime},s^{\prime\prime}\notin S_{i_{j}}$ for any $1\leq j\leq n$;
@@
The existence of $s^{\prime},s^{\prime\prime}\in S_{F}$ as in (i) follows from 5.12(ii) because ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective by assumption. For (ii), we let $F_{1},F_{2}\subseteq\overline{F_{3}}$ with $F=F_{1}$, $S_{F_{2}}=\{s_{j},s^{\prime},s^{\prime\prime}\}$ and $S_{F_{3}}=\{s^{\prime},s^{\prime\prime}\}$, for a suitably chosen $1\leq j\leq n$ so that such an $F_{2}$ exists. In case (a), any choice of $j$ works. In case (b), we pick $j$ such that $\left|S_{i_{j}}\right|>3$. In case (c), we pick $j$ such that $s^{\prime},s^{\prime\prime}\notin S_{i_{j}}$. Either way, as ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi)\neq{\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{2}}}(\chi^{\prime})$, it follows from 5.6(i) and 5.11(iii) that we only need to show that ${\operatorname{Res}}^{{\mathcal{H}_{\operatorname{aff}}}}_{{\mathcal{H}}_{F_{3}}}(\chi)$ is not projective, which itself holds by 5.12(ii). ∎
We now strengthen further our assumptions on $\chi$, $\chi^{\prime}$ and the root system. We write $\Phi=\sqcup_{i=1}^{r}\Phi_{i}$ as before and assume that $S_{1}=\{s,s^{\prime},s^{\prime\prime}\}\subseteq S_{\xi}$, where $\chi(T_{\hat{s^{\prime}}})\neq\chi(T_{\hat{s^{\prime\prime}}})$ and $s^{\prime},s^{\prime\prime}$ are adjacent in the affine Dynkin diagram of $\Phi$. Furthermore, we assume that $s\in S_{1}$ is the unique simple affine reflection for which $\chi(T_{\hat{s}})\neq\chi^{\prime}(T_{\hat{s}})$. By 5.13 and 5.14(ii), we have $f_{F}\sim 0$ for every $F\subseteq\overline{C}$ unless these assumptions hold. With these strengthened assumptions, we now finally show that $f_{F}\sim 0$ for every $F\subseteq\overline{C}$ unless the hypotheses of 5.8(i) are satisfied.
@@
Let $\chi$, $\chi^{\prime}$, $\underline{f}$ be given as in 5.14, and suppose that $\Phi$, $\chi$ and $\chi^{\prime}$ satisfy the assumptions above. Let $F\subseteq\overline{C}$ be any face with $S_{F}\cap S_{1}=\{s^{\prime},s^{\prime\prime}\}$.